The mass, in SI units, of an object that measures 30 cm x 20 cm x 0,1 m and has a density of 0,3kg/m3 is 

A. 0,0018 kg

B. 2g

C.18kg

D. 15kg

E. 20kg


1
Expert's answer
2021-01-19T04:35:22-0500

The volume of the object is:


"V=a*b*c=0.3m*0.2m*0.1m=0.006m\neg3"

The mass of the object is


"m=\rho*V=0.006*0.3=0.0018kg"

The ansver is

A 0.0018 kg